# Posterior Inference with Diffusion Models for High-dimensional Black-box Optimization

### v1

@@ Abstract
Optimizing high-dimensional and complex black-box functions is crucial in numerous scientific applications. While Bayesian optimization (BO) is a powerful method for sample-efficient optimization, it struggles with the curse of dimensionality and scaling to thousands of evaluations. Recently, leveraging generative models to solve black-box optimization problems has emerged as a promising framework. However, those methods often underperform compared to BO methods due to limited expressivity and difficulty of uncertainty estimation in high-dimensional spaces. To overcome these issues, we introduce DiBO, a novel framework for solving high-dimensional black-box optimization problems. Our method iterates two stages. First, we train a diffusion model to capture the data distribution and an ensemble of proxies to predict function values with uncertainty quantification. Second, we cast the candidate selection as a posterior inference problem to balance exploration and exploitation in high-dimensional spaces. Concretely, we fine-tune diffusion models to amortize posterior inference. Extensive experiments demonstrate that our method outperforms state-of-the-art baselines across various synthetic and real-world black-box optimization tasks. Our code is publicly available here.
@@ 1 Introduction
Optimizing high-dimensional and complex black-box functions is crucial in various scientific and engineering applications, including hyperparameter optimization (Snoek et al., 2012; Turner et al., 2021), chemical engineering (Hernández-Lobato et al., 2017), drug discovery (Negoescu et al., 2011), and control systems (Candelieri et al., 2018).
Bayesian optimization (BO) (Brochu et al., 2010; Shahriari et al., 2015) is a powerful method for solving black-box optimization. BO constructs a surrogate model from observed data and finds an input that maximizes the acquisition function to query the black-box function. However, it scales poorly to high dimensions due to the curse of dimensionality (Kandasamy et al., 2015; Letham et al., 2020) and struggles scaling to thousands of evaluations (Wang et al., 2018; Eriksson et al., 2019). To address these challenges, several approaches have been suggested to scale up BO for high-dimensional optimization problems. Some works propose a mapping from high-dimensional space into low-dimensional subspace (Nayebi et al., 2019; Letham et al., 2020) or assume additive structures of the target function (Duvenaud et al., 2011; Rolland et al., 2018) to perform optimization in low-dimensional spaces. However, these methods often rely on unrealistic assumptions.
@@
Recently, generative model-based approaches have shown promising results in black-box optimization (Brookes et al., 2019; Kumar & Levine, 2020; Wu et al., 2024). These methods utilize an inverse mapping from function values to the input domain and propose candidates via sampling from the trained model conditioned on a high score. While they alleviate aforementioned issues in BO by converting optimization problem as sampling (Janner et al., 2022), their performance degrades in higher dimensions due to the limited expressivity of underlying models (Brookes et al., 2019; Kumar & Levine, 2020) or the difficulty of uncertainty estimation in high-dimensional spaces (Wu et al., 2024).
@@
Our method iterates two stages. First, we train a diffusion model to effectively capture high-dimensional data distribution and an ensemble of proxies to predict function values with uncertainty quantification. During training, we adopt a reweighted training scheme proposed in prior generative model-based approaches to focus on high-scoring data points (Kumar & Levine, 2020; Krishnamoorthy et al., 2023). Second, we sample candidates from the posterior distribution. As the sampling from the posterior distribution is intractable, we train an amortized sampler to generate unbiased samples from the posterior by fine-tuning the diffusion model. While the posterior distribution remains highly non-convex, we can capture such complex and multi-modal distribution by exploiting the expressivity of the diffusion models. By repeating these two stages, we progressively get close to the high-scoring regions of the target function.

### v2

@@ Abstract
Optimizing high-dimensional and complex black-box functions is crucial in numerous scientific applications. While Bayesian optimization (BO) is a powerful method for sample-efficient optimization, it struggles with the curse of dimensionality and scaling to thousands of evaluations. Recently, leveraging generative models to solve black-box optimization problems has emerged as a promising framework. However, those methods often underperform compared to BO methods due to limited expressivity and difficulty of uncertainty estimation in high-dimensional spaces. To overcome these issues, we introduce DiBO, a novel framework for solving high-dimensional black-box optimization problems. Our method iterates two stages. First, we train a diffusion model to capture the data distribution and deep ensembles to predict function values with uncertainty quantification. Second, we cast the candidate selection as a posterior inference problem to balance exploration and exploitation in high-dimensional spaces. Concretely, we fine-tune diffusion models to amortize posterior inference. Extensive experiments demonstrate that our method outperforms state-of-the-art baselines across synthetic and real-world tasks. Our code is publicly available here.
@@ 1 Introduction
Optimizing high-dimensional and complex black-box functions is crucial in various scientific and engineering applications, including hyperparameter optimization (Šehić et al., 2022), material discovery (Hernández-Lobato et al., 2017), chemical design (Gómez-Bombarelli et al., 2018; Griffiths et al., 2023), and control systems (Candelieri et al., 2018).
Bayesian optimization (BO) (Kushner, 1964; Garnett, 2023) is a powerful method for solving black-box optimization. BO constructs a surrogate model from observed data and finds an input that maximizes the acquisition function to query the black-box function. However, it scales poorly to high dimensions due to the curse of dimensionality (Wang et al., 2016) and struggles scaling to thousands of evaluations (Wang et al., 2018; Eriksson et al., 2019). To address these challenges, several approaches have been suggested to scale up BO for high-dimensional optimization problems. Some works propose a mapping from high-dimensional space into low-dimensional subspace (Gómez-Bombarelli et al., 2018; Maus et al., 2022; Nayebi et al., 2019; Letham et al., 2020) or assume additive structures of the target function (Duvenaud et al., 2011; Rolland et al., 2018) to perform optimization in low-dimensional spaces. However, these methods often rely on unrealistic assumptions.
@@
Recently, generative model-based approaches have shown promising results in black-box optimization (Brookes et al., 2019; Kumar & Levine, 2020; Wu et al., 2024). These methods utilize an inverse mapping from function values to the input domain and propose candidates via sampling from the trained model conditioned on a high score. While they alleviate aforementioned issues in BO by converting optimization as sampling (Janner et al., 2022), their performance degrades in higher dimensions due to the limited expressivity of underlying models (Brookes et al., 2019; Kumar & Levine, 2020) or the difficulty of uncertainty estimation in high-dimensional spaces (Wu et al., 2024).
@@
Our method iterates two stages. First, we train a diffusion model to effectively capture high-dimensional data distribution and an ensemble of proxies to predict function values with uncertainty quantification. During training, we adopt a reweighted training scheme proposed in prior generative model-based approaches to focus on high-scoring data points (Kumar & Levine, 2020; Tripp et al., 2020). Second, we sample candidates from the posterior distribution. As the sampling from the posterior distribution is intractable, we train an amortized sampler to generate unbiased samples from the posterior by fine-tuning the diffusion model. While the posterior distribution remains highly non-convex, we can capture such complex and multi-modal distribution by exploiting the expressivity of the diffusion models. By repeating these two stages, we progressively get close to the high-scoring regions of the target function.
@@ 7 Conclusion
Acknowledgement
This work is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (No. RS-2025-00563763).